OpenStudy (blu_seaa):

Solve the equation. If necessary, round your answer to the nearest tenth. x2 = 324 A. –18, 18 B. 18 C. –162, 162 D. 162

OpenStudy (welshfella):

its best to write x squared as x^2

OpenStudy (blu_seaa):

That's how the site put it.

OpenStudy (welshfella):

the answer is +/- squared root of 324

OpenStudy (welshfella):

hmm - poor site!

OpenStudy (luckyness):

Okay so when you have x2 it means x*2 So we have x*2=324 Now we have to find out what times 2 can = to 324. Simplify both sides of the equation. 2x=324 Step 2: Divide both sides by 2. \[\frac{ 2x }{ 2 }= \frac{ 324 }{ 5 }\] \[x=162\]

OpenStudy (luckyness):

Answer: D

OpenStudy (blu_seaa):

Thank you.

OpenStudy (luckyness):

No problem :)

OpenStudy (welshfella):

- thats an example of the difficulty that can be cuased:- x2 can be interpreted as ' 2 times x' or 'x squared'.